This loads a font easier to read for people with dyslexia.
This renders the document in high contrast mode.
This renders the document as white on black
This can help those with trouble processing rapid screen movements.

CASS News

Mopra Telescope -- Operations update 13 March 2013


There have been a number of delays to the plan given in the last update, caused in part by wet weather and contractor availability, but the clean-up of the rubble from the site has now been completed. We have been advised to have the equipment in the screened room professionally cleaned before switching on power. The concern is that any ash that may have settled on equipment can absorb moisture from the air and become quite corrosive. The plan is to make a start on this work this week, with the task likely to extend to the end of next week or possibly early the week after.

We had undertaken photogrammetry of the Mopra dish and sub-reflector last October, and so we took another round of photogrammetry last month to see whether the heat of the fire had caused any damage to the surface: the good news is that the two results are consistent with each other and there appear to be no ill-effects due to the bushfire.

Mains power has been restored externally to the site, and our electricians have now installed a temporary switchboard on the control building to enable testing of the equipment. This naturally has to wait until the cleaning of the equipment in the screened room and on the antenna has been completed, so it is likely to be the last week of March before we are able to do the first tests. The current plan is to do the initial tests without cooling the receivers, to confirm that the telescope and MOPS are functioning, and to cool the receivers once the on-site generator has been tested and also proven to be functioning reliably.

Added by Phil Edwards on 2013-03-13

More in General category
More news